OpenStudy (anonymous):

2. For each of the following questions, choose the correct answer. Elena has bought 320 tropical fish for a museum display. She bought 7 times as many triggerfish as parrotfish. How many of each type of fish did she buy? (Points : 1) 280 parrotfish, 40 triggerfish 20 parrotfish, 140 triggerfish 120 parrotfish, 200 triggerfish 40 parrotfish, 280 triggerfish

OpenStudy (anonymous):

What do you know? You know there are parrot fish and trigger fish and when you add them together you have 320 tropical fish. So right away you can answer two is no good. You know that there are 7 times as many trigger fish as parrot fish. So right away you know the answer will have more trigger fish then parrot fish. So you can see that answer one is no good. So now you are left with answer 3 and answer 4. Since there are 7 times as many trigger fish as parrot fish, then there must be 7 trigger fish for one parrot fish. So looking at the last two answers, if we multiply the number of parrot fish by 7 we will know the number of trigger fish. Checking answer 3, we see that 7 times 120 is NOT 200. So that answer is no good. Checking answer 4, we see that 7 times 40 is 280. That answer looks good since 40 + 280 = 320 tropical fish.

OpenStudy (anonymous):

Writing an equation we let P= parrot fish and T= trigger fish. So P + T = 320 fish Since there are 7 times as many trigger fish as parrot fish, wh have this: 7P= T Substituting into the first equation we have this: 7P+P=320 fish 8P=320 fish P=40 So there are 40 parrot fish. Substituting in the second equation we have this: 7*40= 280 So there are 280 trigger fish.